# Local Loss Optimization in the Infinite Width: Stable Parameterization of Predictive Coding Networks and Target Propagation

### v1

@@ B.1.1 Linear Network
In Figure S.1, we measure the similarity of the inference vector with BP, GNT, and the analytical solution. With fewer inference iterations, the model behaves more like BP; however, as the the number of iterations increases, the model converges toward the analytical solution. Furthermore, as the middle layer width $M_{l}$ increases, the gap between GNT and BP decreases. Figure S.2 further demonstrates that reducing the output dimension $M_{L}$ brings the model closer to BP. However, increasing $M_{L}$ moves the model away from BP, though this divergence is more gradual as $\bar{\gamma}_{L}$ approaches zero.
B.1.2 Additional Experiments on $\mu$ transfer for PC
@@
Figure S.4 : Without F-ini, $\mu$P with $\bar{\gamma}_{L}=-1$ transfers the learning rates across width also in MLP. We trained 3-layer MLP on FashionMNIST without F-ini.
@@
In the main text, we mainlyu used MSE loss. However, this can be replaced with cross-entropy (CE) loss. As demonstrated in Figure S.5, $\mu$P for PC also transfers the learning rate across widths when using cross-entropy loss.
@@
Figure S.8 : The results of $\mu$P for PC are independent of the number of training samples. We train 3-layer MLP on FashionMNIST with full training samples. The stability of $\mu$P holds even with all training samples.
We reduced the number of training samples in most of the graphs for $\mu$Transfer. By reducing the number of training samples, finite-width models are known to behave more similarly to infinite-width models, as has often been seen in papers examining the theoretical aspects of feature learning [Geiger et al., 2020, Ishikawa and Karakida, 2024]. However, even when training on the full dataset, $\mu$P remains stable across widths, as shown in Figure S.8.
@@
Figure S.9 : When evaluating the loss after inference, only $\mu$P with $\bar{\gamma}_{L}=-1$ satisfies the empirical rule of “wider is better” Regardless of whether F-ini is applied, $\mu$P with $\bar{\gamma}_{L}=-1$ consistently reduces the loss during inference with stability. We traine 3-layer MLP on FashionMNIST.
@@
As shown in Figure S.11, the shift in the optimal learning rate at $\gamma_{L}=0$ with $M^{\prime}=128$ becomes more evident as the number of inference iterations increases. This is likely because with more iterations, the dynamics of inference play a more critical role in weight updates. In summary, to achieve stable $\mu$Transfer independent of the base width and the number of inference iterations, we should use $\mu$P with $\gamma_{L}=-1$.
Figure S.10 : When the base width $M^{\prime}$ is large, $\mu$P with $\bar{\gamma}_{L}=0$ tends to fail with $\mu$transfer. We train 3-layer MLP on FashionMNIST. This suggests that $\mu$P with $\gamma_{L}=-1$ should be used when setting the base width, even with F-ini. The inference is performed for 100 iterations.
@@
Figure S.12 : $\mu$P for PC also transfers learning rates across widths in synchronous inference. $\mu$P for PC can also be applied in synchronous inference. Note that when the base width is set to 1, as in SI, the learning rate does not transfer in $\mu$P with $\bar{\gamma}_{L}=0$.
Algorithm 1 PC Algorithm (Simultaneous or Sequential inference)
1:for $s=1$ to $n$ do
3:     for $l=L-1$ to $1$ do
4:         $e_{l,s}\leftarrow u_{l,s}-W_{l}\phi_{l-1}(u_{l-1,s})$
@@
Refer to caption Refer to caption
Figure S.13 : $\mu$P scales better than SP and NTK. We trained a CNN by PC with F-ini and a VGG5 on the full FashionMNIST dataset. The "wider is better" principle holds for $\mu$P.
Figure S.13 presents the results of the same experiment shown in Figure 3 (Right), but with the CIFAR-10 dataset and the VGG5 model. It is evident that even with CIFAR-10 and VGG5, $\mu$P achieves higher accuracy compared to SP and NTK.
B.2 Target Propagation
B.2.1 Additional Experiments on $\mu$Transfer for PC
@@
As discussed in Section A.5, stable parametrization is crucial not only for feedforward but also for feedback networks. We verified this with $\mu$P, as shown in Figures S.14 and S.15.
@@
Meulemans et al. [2020] proposes the difference reconstruction loss (DRL) for constructing feedback networks. In Figure S.16, we empirically confirm that our $\mu$P works effectively with DRL when training an MLP on FashionMNIST.
@@
Appendix C Experimental Settings
@@
  • CNN: We trained 3-layer CNN with Tanh activation. The models consist of a two-layer convolutional layers and a linear layer. We trained with different hidden widths where the width is proportional to the input dimension of the output layer. (For example, when the width is set to 4, the input dimension of the final layer is 512.) Max pooling is applied after the activation function.
  • VGG5: We trained a VGG-like model consisting of 4 convolutional blocks and 3 linear blocks, based on the structure described in [Pinchetti et al., 2024]. When the width is set to 8, it matches the VGG5 model in Pinchetti et al. [2024], with the channel sizes being [128, 256, 512, 512].
Dataset and batch size.
We used FahionMNIST and CIFAR-10 datasets without applying any data augmentation. The settings for batch size and training samples were as follows:
@@
  • TP In Figures 6 and 7, we trained 3-layer MLP using 1024 training samples. Note that in Figure S.17 in the Appendix, FashionMNIST and CIFAR-10 were trained using the full datasets.

### v2

@@ D.1.1 Linear Network
In Figure S.1, we measure the similarity of the inference vector with BP, GNT, and the analytical solution. With fewer inference iterations, the model behaves more like BP; however, as the number of iterations increases, the model converges toward the analytical solution. Furthermore, as the middle layer width $M_{l}$ increases, the gap between GNT and BP decreases. Figure S.2 further demonstrates that reducing the output dimension $M_{L}$ brings the model closer to BP. However, increasing $M_{L}$ moves the model away from BP, though this divergence is more gradual as $\bar{\gamma}_{L}$ approaches zero.
D.1.2 Additional Experiments on $\mu$ transfer for PC
@@
Figure S.4 : Without F-ini, $\mu$P with $\bar{\gamma}_{L}=-1$ transfers the learning rates across widths also in MLP. We trained a 3-layer MLP on FashionMNIST without F-ini.
@@
In the main text, we mainly used mean squared error (MSE) loss. However, this can be replaced with cross-entropy (CE) loss. As demonstrated in Figure S.5, $\mu$P for PC also transfers the learning rate across widths when using cross-entropy loss.
@@
Figure S.8 : The results of $\mu$P for PC are independent of the number of training samples. We train a 3-layer MLP on FashionMNIST with full training samples. The stability of $\mu$P holds even with all training samples.
We reduced the number of training samples in most of the graphs for $\mu$Transfer. By reducing the number of training samples, finite-width models are known to behave more similarly to infinite-width models, as has often been seen in papers examining the theoretical aspects of feature learning (Geiger et al., 2020; Ishikawa & Karakida, 2024). However, even when training on the full dataset, $\mu$P remains stable across widths, as shown in Figure S.8.
@@
Figure S.9 : When evaluating the loss after inference, only $\mu$P with $\bar{\gamma}_{L}=-1$ satisfies the empirical rule of “wider is better” Regardless of whether F-ini is applied, $\mu$P with $\bar{\gamma}_{L}=-1$ consistently reduces the loss during inference with stability. We trained a 3-layer MLP on FashionMNIST.
@@
As shown in Figure S.11, the shift in the optimal learning rate at $\gamma_{L}=0$ with $M^{\prime}=128$ becomes more evident as the number of inference iterations increases. This is likely because, with more iterations, the dynamics of inference play a more critical role in weight updates. In summary, to achieve stable $\mu$Transfer independent of the base width and the number of inference iterations, we should use $\mu$P with $\gamma_{L}=-1$.
Figure S.10 : When the base width $M^{\prime}$ is large, $\mu$P with $\bar{\gamma}_{L}=0$ tends to fail with $\mu$transfer. We train a 3-layer MLP on FashionMNIST. This suggests that $\mu$P with $\gamma_{L}=-1$ should be used when setting the base width, even with F-ini. The inference is performed for 100 iterations.
@@
Figure S.12 : $\mu$P for PC also transfers learning rates across widths in synchronous inference. $\mu$P for PC can also be applied in synchronous inference. Note that when the base width is set to 128, as in SI, the learning rate does not transfer in $\mu$P with $\bar{\gamma}_{L}=0$.
@@
Figure S.13 : $\mu$P scales better than SP and NTK. We trained a CNN by PC with F-ini on CIFAR10/CIFAR100 and a VGG5 on the full FashionMNIST dataset. The "wider is better" principle holds for $\mu$P.
Figure S.13 presents the results of the same experiment shown in Figure 3 (Right), but with the CIFAR-10/CIFAR=100 dataset and the VGG5 model. It is evident that even with CIFAR-10, CIFAR-100 and VGG5, $\mu$P achieves higher accuracy compared to SP and NTK.
D.2 Target Propagation
D.2.1 Additional Experiments on $\mu$Transfer for TP
Figure S.14 : In Target Propagation, using $\mu$P ensures that $\Delta h_{l}$ remains consistent across widths. This figure shows the RMS norm of $\Delta h_{l}$ during training. For SP, $\Delta h_{l}$ in the input layer diminishes as the width increases, while $\Delta h_{l}$ in the output layer diverges with increasing width. Consequently, the training dynamics become unstable. In contrast, with $\mu$P, $\Delta h_{l}$ remains consistent across different widths in both the input and output layers.
@@ D.2.1 Additional Experiments on $\mu$Transfer for TP
Temporal change of activation
In Figure S.14, we observed $\Delta h$ during the training of an MLP on FashionMNIST. SP exhibits a dependency of $\Delta h$ on width, whereas $\mu$P demonstrates consistent behavior, independent of width.
@@
As discussed in Section C.2, stable parameterization is crucial not only for feedforward but also for feedback networks. We verified this with $\mu$P, as shown in Figures S.15 and S.16.
@@
Meulemans et al. (2020) proposes the difference reconstruction loss (DRL) for constructing feedback networks. In Figure S.17, we empirically confirm that our $\mu$P works effectively with DRL when training an MLP on FashionMNIST.
@@
Appendix E Experimental Settings
@@
  • CNN: We trained a 3-layer CNN with Tanh activation. The models consist of two-layer convolutional layers and a linear layer. We trained with different hidden widths where the width is proportional to the input dimension of the output layer. (For example, when the width is set to 4, the input dimension of the final layer is 512.) Max pooling is applied after the activation function.
  • VGG5: We trained a VGG-like model consisting of 4 convolutional blocks and 3 linear blocks, based on the structure described in (Pinchetti et al., 2024). When the width is set to 8, it matches the VGG5 model in Pinchetti et al. (2024), with the channel sizes being [128, 256, 512, 512].
Dataset and batch size
We used FashionMNIST and CIFAR-10 datasets without applying any data augmentation. The settings for batch size and training samples were as follows:
@@
  • TP In Figures S.14 and 6, we trained a 3-layer MLP using 1024 training samples. Note that in Figure S.18 in the Appendix, FashionMNIST, and CIFAR-10 were trained using the full datasets. For the activation function of feedback networks, the same activation function as one used in the forward pass is utilized (i.e., $\psi=\phi$).